# eprint Nisho-2-2025

### v3

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B_{0}$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface. In contrast, if the cylinder is composed of a material with small conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder. Within the QCD axion model, the power $P$ is estimated as $P\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. The detection of dark matter axions in our method may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
$I$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
@@
In the observation, the power $P(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $P_{n}=T\delta\omega/2\pi$ with $\delta\omega=10^{-6}m_{a}$,
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{n}}\simeq 8.6\times 10^{-7}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{B}{15\text{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (4)
The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=100$seconds. In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $P(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/P_{n}\simeq 1.3\times 10^{-3}$.
Although the details of power detection are beyond the scope of this paper, we expect the ratio to be sufficiently large for the power ( $P(\sigma=10^{-3},\text{eV})$ ) to become detectable in the near future. Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of the so-called dark photon. The electromagnetic coupling of the dark photon produces an effect similar to that of the axion.
@@
$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big),$ (27)
@@
To achieve a signal-to-noise ratio greater than 1, the power $P$ must be amplified by approximately three orders of magnitude. The method of amplification and its technical implementation are beyond the scope of this paper. However, once such amplification is realized, the dark matter axion can be detected using our proposed approach.
@@
Summarized, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$. With such conductivity, the axion induces large electric power generated in the bulk of the cylinder and causes a high signal-to-noise ratio. Therefore, the method demonstrates the feasibility of detecting dark matter axions under realistic experimental conditions. It is also effective for the search of dark photon.

### v4

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B_{0}$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface. In contrast, if the cylinder is composed of a material with small conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder. Within the QCD axion model, the power $P$ is estimated as $P\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. The signal of the axion can be detected by observing electric current in parallel LC circuit with quality factor $Q\geq 2$. The detection of dark matter axions in our method may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
$I(\sigma=10^{-3}\text{eV})$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (1)
@@
In the observation, the electric $I(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ where the resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder and $\delta\omega=10^{-6}m_{a}$,
$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 4.2\times 10^{-4}g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\mbox{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{B}{15\text{T}}\Big)\Big(\frac{10^{-4}\text{eV}}{m_{a}}\Big)^{1/2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (4)
The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=100$seconds. In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $I(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 6.7\times 10^{-1}$.
The detection of the current is observed using parallel LC circuit in which the current is amplified with resonance by tuning parameters of the circuit. Even when we have small enhancement factor $Q=10$, the signal-noise ratio becomes larger than $1$; $QI(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 4.6$. Consequently, the signal of the axion can be detected by our method.
Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of the so-called dark photon. The electromagnetic coupling of the dark photon produces an effect similar to that of the axion.
@@
$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$ (27)
@@
These powers can be detected using a parallel LC circuit, in which electric current in coil is measured. By tuning the parameters of the LC circuit, the electric current at resonance is enhanced by a factor $Q=R_{c}/m_{a}L_{c}$ where $L_{c}$ denotes inductance of the inductor and resistance $R_{c}=R$ of the cylinder. The cylinder forms the resistor of the circuit. The resonance condition is given by $m_{a}=1/\sqrt{C_{c}L_{c}}$ where $C_{c}$ is capacitance of the capacitor.
In order to observe the electric current $I(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-14}$A, we compare it with the thermal noise $I_{n}$,
$I_{n}(\sigma=10^{-3}\text{eV})=\sqrt{\frac{2T\delta\omega}{\pi R_{c}}}\simeq 6.6\times 10^{-11}\text{A}\Big(\frac{T}{20\text{mK}}\Big)^{1/2}\Big(\frac{100\text{cm}}{L}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)^{1/2}\Big(\frac{m_{a}}{10^{-4}\text{eV}}\Big)^{1/2}$ (32)
with the resistance $R_{c}=L/(\sigma\pi R^{2})\simeq 1.44(L/100\text{cm})(6\text{cm}/R^{2})\simeq 540\Omega$. Then, a signal-to-noise ratio at resonance is given by
$\frac{QI(\sigma=10^{-3}\text{eV})}{I_{n}({\sigma=10^{-3}\text{eV})}}$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq$ (33)
$0.66Q$ $g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{B_{0}}{15\mbox{T}}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)^{1/2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}$
Consequently, to achieve the signal-to-noise ratio greater than 1, it is sufficient to have $Q\geq 2$. Of course, it is favorable to have much large $Q$ value for the axion detection. Once such a quality factor $Q$ is realized, the dark matter axion can be detected using our proposed approach.
@@
Summarized, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$. With such conductivity, the axion induces large electric current generated in the bulk of the cylinder. It causes a high signal-to-noise ratio of the current in parallel LC circuit even with small $Q\geq 2$. Therefore, the method demonstrates the feasibility of detecting dark matter axions under realistic experimental conditions. It is also effective for the search of dark photon.